Can you please show me step by step how to do: 2b + 4(b - 1)= 5(b-1)?
This is my first time doing this(freshman).
Thank You

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

2b+4(b-1)=5(b-1)
there are at least two ways to do this. one easy and one easier
we can subtract 4(b-1) from both sides
2b=1(b-1)
2b=b-1
subtract b from both sides
b=-1
another way
2b+4(b-1)=5(b-1)
expand
2b+4b-4=5b-5
combine like terms
6b-4=5b-5
subtract 5b from both sides
b-4=-5
add 4 to both sides
b=-1

Answer by solver91311(24713)





Multiply 4 times b and 4 times -1:



Multiply 5 times b and 5 times -1:



Add -5b to both sides:



And collect like terms:



Add 4 to both sides:





Check:







Yep, every time.

John